The greater of the two consecutive even integers is 20 more than twice the lesser. find the integers.. I really dont even know where to start it confuses me...
Answer by stanbon(75887)

The greater of the two consecutive even integers is 20 more than twice the lesser. find the integers..
-------
1st: 2x
2nd: 2x+2
-----
Equation:
2x+2 = 2(2x)+20
2x + 2 = 4x + 20
---
Solve to get:
1st: 2x = -18
2nd: 2x+2 = -16